# GW231123: a Binary Black Hole Merger with Total Mass 190-265

### v2

@@ Abstract
On 2023 November 23 the two LIGO observatories both detected GW231123, a gravitational-wave signal consistent with the merger of two black holes with masses and (90% credible intervals), at luminosity distance 0.7–4.1 Gpc and redshift of , and a network signal-to-noise ratio of 22.5. Both black holes exhibit high spins, and respectively. A massive black hole remnant is supported by an independent ringdown analysis. Some properties of GW231123 are subject to large systematic uncertainties, as indicated by differences in inferred parameters between signal models. The primary black hole lies within or above the theorized mass gap where black holes between 60–130 should be rare due to pair instability mechanisms, while the secondary spans the gap. The observation of GW231123 therefore suggests the formation of black holes from channels beyond standard stellar collapse, and that intermediate-mass black holes of mass 200 form through gravitational-wave driven mergers.
]Compiled: August 11, 2025
@@ 1 Introduction
From 2015 to 2020 the LIGO-Virgo-KAGRA Collaboration identified 69 gravitational-wave signals from binary black hole mergers with false alarm rates below one per year (Aasi et al., 2015; Acernese et al., 2015; Abbott et al., 2020c; Akutsu et al., 2020; Abbott et al., 2023a). Of these, the most massive was the source of GW190521, with a merger remnant of 140 (Abbott et al., 2020d, e). The small number of observable cycles of GW190521 limits our ability to accurately infer the source’s properties, and subsequent studies have proposed a wide range of alternative interpretations, including highly eccentric orbits, dynamical capture scenarios, exotic object mergers, and cosmic string collapse (Gayathri et al., 2022; Romero-Shaw et al., 2020a; Gamba et al., 2023; Calderón Bustillo et al., 2021b; Aurrekoetxea et al., 2024). Here we present a yet more challenging signal: GW231123_135430 (hereafter referred to as GW231123), confidently observed through a coincident detection in both the LIGO Hanford and Livingston detectors during the first part of their fourth observing run, O4a (2023 May 24 to 2024 January 16). The combination of data from the two observatories was essential in making a confident detection.
@@
GW231123 consists of 5 cycles over a frequency range of 30–80 Hz. We interpret GW231123 as a binary-black-hole merger and infer a total mass between 190 and 265 and high component black-hole spins (0.9 and 0.8). While a few gravitational-wave candidates have been observed with similarly high total masses (Abbott et al., 2023a; Wadekar et al., 2023), none have false alarm rates less than 1 per year; in addition, GW231123 has both a large signal-to-noise ratio and high statistical significance. Such high masses and spins pose a challenge to our most accurate waveform models, leading to larger uncertainties in the black-hole masses, and the binary orientation and distance, than any previous signal of comparable strength.
@@
Pair instability supernovae and pulsational pair instability supernovae are expected to preclude stellar collapse to black holes with masses –130 (Farmer et al., 2019, 2020; Woosley & Heger, 2021; Hendriks et al., 2023), and the majority of the astrophysical population of black holes inferred from gravitational-wave catalogs lies below this gap (Abbott et al., 2023b). The large measurement uncertainties in the source of GW231123 mean that the primary black hole may be within or beyond the mass gap, while the range of possible values for the secondary mass spans the entire gap. It is also possible that the two black-hole masses lie on either side of the gap. The scenarios with the highest probability require a formation channel that populates the mass gap, such as prior stellar (e.g., Di Carlo et al., 2020; Renzo et al., 2020a; Kremer et al., 2020) or black-hole (Gerosa & Fishbach, 2021) mergers. The high inferred spins do not exclude these interpretations, although their values may be higher than the inferred spins of the remnant black holes produced in the binary-black-hole mergers observed so far.
@@ 2 Detection Significance
On 2023 November 23, at 13:54:30 UTC, the Advanced LIGO Hanford and Livingston detectors observed the GW transient GW231123. Despite its short duration (0.1 s) and limited bandwidth (Figure 1), the signal was identified in low-latency analyses with high statistical significance, reported in terms of inverse false-alarm rate (IFAR); see Table. 1. It was first detected by PyCBC Live, a matched-filter search for compact binaries (Allen, 2005; Usman et al., 2016; Nitz et al., 2017; Dal Canton et al., 2021). It was also reported by coherent WaveBurst (cWB)-BBH, a minimally modelled coherent excess power search (Mishra et al., 2025). The cWB-BBH search uses the WaveScan time–frequency (TF) transformation (Klimenko, 2022) and ranks identified triggers using a machine-learning classifier trained specifically on BBH signals (Mishra et al., 2021, 2022, 2025). In addition, the event was also detected by two model-independent low-latency cWB searches, or burst searches, designed to identify generic GW transients: cWB-2G and cWB-XP. The former is based on the Wilson–Debauchies–Meyer TF transformation (Klimenko et al., 2008, 2016a; Drago et al., 2020), while the latter uses the WaveScan TF transformation. Both apply a machine-learning classifier (XGBoost), trained on generic white-noise-bursts to rank identified triggers (Szczepańczyk et al., 2023). For further details on model-independent searches, see Abac et al. (2025a).
Subsequent offline or archival reanalyses using improved background estimation and data quality information further increased the event’s statistical significance in both the PyCBC and cWB pipelines (Table 1). Furthermore, two additional matched-filter searches, GstLAL (Messick et al., 2017; Sachdev et al., 2019; Hanna et al., 2020; Cannon et al., 2021; Sakon et al., 2024; Ewing et al., 2024; Tsukada et al., 2023; Joshi et al., 2025) and MBTA (Adams et al., 2016; Aubin et al., 2021; Alléné et al., 2025), which did not detect this event with significant confidence in low-latency (IFAR higher than 1 year), recovered it in their offline analyses. These searches differ from PyCBC (Davies et al., 2020; Chandra et al., 2021b; Davis et al., 2022; Kumar & Dent, 2024) in their implementation and use of signal–noise discriminators. GstLAL’s enhanced significance (the higher IFAR in Table. 1) is primarily driven by a higher mass extension of the search with specific settings to compute the background for such higher mass mergers accurately. More details are provided in Abac et al. (2025b). These changes improved the signal and noise models in this part of the parameter space in general, leading to a better recovery of this high mass signal. Additionally, cWB-GMM, an entirely offline model-independent search, uses Gaussian Mixture Models (Gayathri et al., 2020; Lopez et al., 2022; Smith et al., 2024) to rerank the triggers identified by cWB-2G.
@@
The differences in the IFARs reported by the offline search pipelines—despite broadly consistent signal-to-noise ratios—primarily reflect differences in their ability to separate GW231123-like signals from background noise in a comparable parameter range. Similar discrepancies have been observed previously, particularly between matched-filter and minimally-modelled searches, when searching for non-eccentric intermediate-mass black hole (IMBH) binaries (Calderón Bustillo et al., 2018; Chandra et al., 2020; Abbott et al., 2020d; Chandra et al., 2021a; Szczepańczyk et al., 2021). These differences arise not only from how effectively each search separates signals from glitches, but also from the differing approaches used to estimate the noise background.
To assess whether the observed variation in statistical significance across pipelines is consistent with expectations, we conducted a dedicated injection campaign. Using the NRSur7dq4 (NRSur) waveform model (Varma et al., 2019), we simulated 8000 non-eccentric BBH signals with intrinsic parameters consistent with those inferred for GW231123 (Section 4). We sampled the sky positions and binary orientations isotropically and drew redshifts uniformly in comoving volume up to , assuming a flat CDM cosmology (Ade et al., 2016). We added these simulated signals uniformly over several days around the event and re-ran our offline search pipelines using the same configuration as applied to the real data.
@@
Given the reported IFAR for GW231123 across the different searches and the results of the injection study, we consider GW231123 to be a confident detection with strong support for an astrophysical origin.
@@ 3 Data quality
The event GW231123 was detected during the first part of the fourth observing run (O4a), a time when the LIGO Hanford and LIGO Livingston detectors were observing with a typical binary neutron star inspiral range of 152 Mpc and 160 Mpc (Capote et al., 2025). The detectors’ data were calibrated in near real-time to produce the online dataset used for low-latency searches (Abbott et al., 2020b; Klimenko et al., 2016b; Tsukada et al., 2023; Ewing et al., 2024; Dal Canton et al., 2021; Chu et al., 2022; Aubin et al., 2021) and parameter estimation (Singer & Price, 2016; Ashton et al., 2019; Pankow et al., 2015). The calibration process subtracts linear spectral features from known instrumental sources, identified through auxiliary witness sensors, and intentionally injected calibration lines used to measure the instruments’ response at various frequencies (Viets et al., 2018a; Sun et al., 2020, 2021).
@@
Following the data-quality procedures established for O4a (Soni et al., 2025), including broadband noise subtraction (Vajente et al., 2020) and data-quality report analysis routines (Davis et al., 2021), detector data surrounding the event were evaluated for signs of non-Gaussian excess power (glitches) within the target time-frequency analysis window using a spectrogram-based glitch-identification tool (Vazsonyi & Davis, 2023). It was determined that glitches were present in each detector around, but not coincident with the event.
@@
From spectrograms, we determined that a glitch was present in the LIGO Hanford data 1.7–1.1 s before the event, in a frequency range between 15–30 Hz. The glitch is possibly related to the LIGO Hanford differential arm control loop (Aasi et al., 2015). This control loop leads to nonstationary noise from the high root-mean-square drive applied to the electrostatic drive actuator. This issue has been fixed in the second part of the fourth observing run (Vajente, 2024). This glitch was close to the event and within the time–frequency window used to infer the source properties, so BayesWave (Cornish & Littenberg, 2015; Cornish et al., 2021; Chatziioannou et al., 2021) was used to model simultaneously the compact binary signal and the glitch (Soni et al., 2025). We removed this non-Gaussianity from the data by subtracting a phenomenological, wavelet-based model of the excess power noise (Hourihane et al., 2022; Ghonge et al., 2024). The glitch-subtracted data successfully passed the validation process, which compares the residual noise to Gaussian noise (Soni et al., 2025; Vazsonyi & Davis, 2023). Additional broadband non-stationary noise was present in the Hanford detector in the hours of data surrounding GW231123, but we found no evidence that this impacted the analysis of GW231123.
@@
In LIGO Livingston data, a glitch was identified 3.0–2.0 s before the event, in a frequency range between 10–20 Hz. Given that LIGO Livingston had recurring low-frequency scattered light glitches (Soni et al., 2025), this glitch was likely caused by scattered light. We determined the time–frequency profile of the glitch to have no measurable effect on the GW231123 analysis, so the analyses from here on use the LIGO Livingston original data and the LIGO Hanford glitch-subtracted data.
@@ 4.1 Methods
We report the properties of GW231123 assuming a non-eccentric BBH merger. We perform a coherent Bayesian analysis (Abbott et al., 2016a) of the LIGO Hanford and LIGO Livingston data around the time of GW231123. We calculate the likelihood using 8 s of data (6 s before and 2 s after the reported merger time of GW231123), and consider frequencies within the range 20–448 Hz. All analyses employ standard priors used in previous analyses (Abbott et al., 2021a, 2024a, 2023a), and we use a Planck 2015 cosmology (Ade et al., 2016). The NRSur analysis employs a reduced prior mass range due to model constraints, mass ratios below 6:1. The other models employ a wider mass prior, mass ratios below 10:1, and no posterior support is found beyond the NRSur analysis. We characterize the detector noise via the median on-source power spectral density (PSD) produced from BayesWave (Cornish & Littenberg, 2015; Littenberg & Cornish, 2015). To sample the posterior distribution, we interface with dynesty (Speagle, 2020) via the bilby library (Ashton et al., 2019; Romero-Shaw et al., 2020b). We use live points and the bilby-implemented acceptance-walk sampling algorithm. We verify that the results remain consistent when the number of live points is increased, as well as when we lower the frequency range to include data between 16–20 Hz.
@@ 4.2 Waveform Systematics
The source properties of GW231123 lie in a challenging region of parameter space for current waveform models, to such an extent that measurements using different models show significant disagreement, with multiple parameters failing to agree within 90% credible intervals. (See Appendix A for examples.) In general, our models are well within our observations’ accuracy requirements, and the level of model disagreement for GW231123 has not been seen in any previous LVK GW observation with moderate SNRs (12). All models show strong support for spins 0.8, and since no theoretical signal model is calibrated to numerical-relativity (NR) waveforms from precessing binaries with spins above 0.8, waveform uncertainties are one possible cause of the measurement differences. Hence, before presenting the source properties, we describe how we quantify waveform-model uncertainties. We do not study in detail the impact of Gaussian noise fluctuations or low-SNR glitches that are difficult to identify and mitigate using the methods presented in Section 3.
@@
We consider five state-of-the-art signal models, NRSur7dq4 (NRSur; Varma et al., 2019), SEOBNRv5PHM (v5PHM; Ramos-Buades et al., 2023a), IMRPhenomTPHM (TPHM; Estellés et al., 2022a), IMRPhenomXPHM (XPHM; Colleoni et al., 2025) and IMRPhenomXO4a (XO4a; Thompson et al., 2024). (More details are given in Appendix A.) The model papers referenced here include studies to assess the accuracy of these models across the BBH parameter space, but here we focus on the likely region of parameter space for this observation; high total mass, , and moderate to high spins.
@@
We quantify the models’ accuracy against NR results, including a set of simulations that extend up to spins of 0.95 (Boyle et al., 2019; Hamilton et al., 2024; Scheel et al., 2025). A standard waveform accuracy measure is the mismatch between two waveforms (Cutler & Flanagan, 1994), where waveform uncertainties will not bias a parameter measurement if the model’s mismatch uncertainty is less than (McWilliams et al., 2010; Baird et al., 2013), where is the SNR and is the chi-square value for degrees of freedom at probability . For single-parameter measurements provides a lower bound (Thompson et al., 2025), so the mismatch criterion for the 90% credible interval at is . Figure 2 reports the distribution of mismatches of each model against 1123 NR waveforms with , all scaled to the redshifted (detector-frame) total mass , at six equally spaced inclinations in from to inclusive. The mismatches are calculated for precessing systems (Schmidt et al., 2015; Harry et al., 2016) following the procedure described in Hamilton et al. (2021), maximising over time shifts, a global phase and template polarisation and optimising over in-plane spin rotations. A subset of the simulations come from the third release of the SXS catalog (Scheel et al., 2025) and contain GW memory, which introduces a constant late-time offset that we handle with a highpass filtering technique (Xu et al., 2024; Valencia et al., 2024; Chen et al., 2024) to mitigate possible artifacts in the Fourier domain. We employ the same PSD for the LIGO Livingston detector as utilised in the coherent Bayesian analysis. NRSur performs better than the other models (by roughly an order of magnitude for low-spin cases), and all other models have comparable accuracy. However, NRSur does not meet the conservative accuracy criterion for all cases, and for spins greater than 0.8 the mismatches are higher in 10% of 98 cases. Even if we apply a less conservative mismatch criterion from the literature (e.g., with for non-eccentric binaries and (Chatziioannou et al., 2017; Scheel et al., 2025) we have 0.0072) there are configurations where NRSur exceeds the criterion (2%). The relative accuracy of models is also not uniform across all cases, e.g., we find cases in which other models show comparable or improved performance relative to NRSur.
@@
To test whether these waveform uncertainties will incur biases, we performed our standard Bayesian parameter estimation analysis on a series of NR injections, as detailed in Appendix A. We observe that, while the five models considered here perform well for most signals, there are configurations where all models may incur biases for massive high-spin signals. We also find that the relative performance of each model can change in the presence of Gaussian noise, although this requires more detailed study in future work. To properly correct for this, we would ideally marginalise over waveform uncertainties or incorporate model accuracy into Bayesian analyses (Read, 2023; Khan, 2024; Hoy et al., 2024; Pompili et al., 2024; Kumar et al., 2025; Mezzasoma et al., 2025). Without access to a model of the waveform-model uncertainties, we follow what has been done previously (Abbott et al., 2016a) and combine the results from multiple models to marginalise over the model uncertainties. In choosing models in addition to NRSur, we note that all other models exhibit a comparable range of mismatches, and no model is clearly preferred in our injection studies, and so we include all five state-of-the-art models. We combine posterior results inferred from NRSur, v5PHM, TPHM, XPHM, XO4a with equal weight and report the combined samples throughout this paper. To illustrate the variation between the combined results and single models, in some figures we also show the NRSur results. In some analyses we expect the choice of model to have little impact, e.g., the detection significance study in Section 2, and in these cases, we use only the NRSur samples.

### v3

@@ Abstract
On 2023 November 23 the two LIGO observatories both detected GW231123, a gravitational-wave signal consistent with the merger of two black holes with masses and (90% credible intervals), at luminosity distance 0.7–4.1 Gpc and redshift of , and a network signal-to-noise ratio of 20.7. Both black holes exhibit high spins, and respectively. A massive black hole remnant is supported by an independent ringdown analysis. Some properties of GW231123 are subject to large systematic uncertainties, as indicated by differences in inferred parameters between signal models. The primary black hole lies within or above the theorized mass gap where black holes between 60–130 should be rare due to pair instability mechanisms, while the secondary spans the gap. The observation of GW231123 therefore suggests the formation of black holes from channels beyond standard stellar collapse, and that intermediate-mass black holes of mass 200 form through gravitational-wave driven mergers.
]Compiled: November 10, 2025
@@ 1 Introduction
From 2015 to 2020 the LIGO-Virgo-KAGRA Collaboration identified 69 gravitational-wave signals from binary black hole mergers with false alarm rates below one per year (LIGOScientific:2014pky; VIRGO:2014yos; abbott2020prospects; 10.1093/ptep/ptaa125; LIGOScientific:2021vkt). Of these, the most massive was the source of GW190521, with a merger remnant of 140 (LIGOScientific:2020iuh; LIGOScientific:2020ufj). The small number of observable cycles of GW190521 limits our ability to accurately infer the source’s properties, and subsequent studies have proposed a wide range of alternative interpretations, including highly eccentric orbits, dynamical capture scenarios, exotic object mergers, and cosmic string collapse (Gayathri:2020coq; Romero-Shaw:2020thy; Gamba:2021gap; CalderonBustillo:2020fyi; Aurrekoetxea:2023vtp). Here we present a yet more challenging signal: GW231123_135430 (hereafter referred to as GW231123), confidently observed through a coincident detection in both the LIGO Hanford and Livingston detectors during the first part of their fourth observing run, O4a (2023 May 24 to 2024 January 16). The combination of data from the two observatories was essential in making a confident detection.
@@
GW231123 consists of 5 cycles over a frequency range of 30–80 Hz, similar to GW190521. We interpret GW231123 as a binary-black-hole merger and infer a total mass between 190 and 265 and high component black-hole spins (0.9 and 0.8). While a few gravitational-wave candidates have been observed with similarly high total masses (LIGOScientific:2021vkt; Wadekar:2023gea), none have false alarm rates less than 1 per year; in addition, GW231123 has both a large signal-to-noise ratio and high statistical significance. Such high masses and spins pose a challenge to our most accurate waveform models, leading to larger uncertainties in the black-hole masses, and the binary orientation and distance, than any previous signal of comparable strength.
@@
PISN and pulsational pair-instability supernovae are expected to preclude stellar collapse to black holes with masses –130 (2019ApJ...887...53F; Farmer:2020xne; Woosley:2021xba; Hendriks:2023yrw), and the majority of the astrophysical population of black holes inferred from gravitational-wave catalogs lies below this gap (KAGRA:2021duu). The large measurement uncertainties in the source of GW231123 mean that the primary black hole may be within or beyond the mass gap, while the secondary mass spans the entire gap within the 90% credible intervals of our analysis. It is also possible that the two black-hole masses lie on either side of the gap. The scenarios with the highest probability require a formation channel that populates the mass gap, such as prior stellar mergers (e.g., DiCarlo:2019fcq; Renzo:2020smh; Kremer:2020wtp), black-hole mergers (for a review, see Gerosa:2021mno and references therein), or accretion in a gaseous environment (e.g., McKernan:2012rf). Though these channels can produce highly spinning black holes (e.g., a characteristic value of for black-hole mergers), the inferred spins may be higher than typical of remnant black holes and those of remnants from mergers previously observed with gravitational waves.
@@ 2 Detection Significance
On 2023 November 23, at 13:54:30 UTC, the Advanced LIGO Hanford and Livingston detectors observed the GW transient GW231123. The Advanced Virgo and KAGRA detectors were not online at this time. Despite its short duration (0.1 s) and limited bandwidth (Figure 1), coherent detection in both detectors allowed the signal to be identified in our analyses with high statistical significance, reported in terms of inverse false-alarm rate (IFAR); see Table 1. Without coincidence in two or more detectors, a high-mass BBH signal like GW231123 would likely have been dismissed as a noise artefact (glitch). It was first detected by PyCBC Live, a matched-filter search for compact binaries (Allen:2004gu; Usman:2015kfa; Nitz:2017svb; DalCanton:2020vpm). It was also reported by coherent WaveBurst (cWB)-BBH, a minimally modelled coherent excess power search (Mishra:2024zzs). The cWB-BBH search uses the WaveScan time–frequency (TF) transformation (Klimenko:2022nji) and ranks identified triggers using a machine-learning classifier trained specifically on BBH signals (Mishra:2021tmu; Mishra:2022ott; Mishra:2024zzs). In addition, the event was also detected by two model-independent low-latency cWB searches, or burst searches, designed to identify generic GW transients: cWB-2G and cWB-XP. The former is based on the Wilson–Debauchies–Meyer TF transformation (Klimenko:2008fu; Klimenko:2015ypf; Drago:2020kic), while the latter uses the WaveScan TF transformation. Both apply a machine-learning classifier (XGBoost), trained on generic white-noise-bursts to rank identified triggers (Szczepanczyk:2022urr). For further details on model-independent searches, see O4a:Allskyshort.
Subsequent offline or archival reanalyses using improved background estimation and data quality information further increased the event’s statistical significance in both the PyCBC and cWB pipelines (Table 1). Furthermore, two additional matched-filter searches, GstLAL (Messick:2016aqy; Sachdev:2019vvd; Hanna:2019ezx; Cannon:2020qnf; Sakon:2022ibh; ewing2023performance; Tsukada:2023edh; Joshi:2025nty) and MBTA (Adams:2015ulm; Aubin:2020goo; Allene:2025saz), which did not detect this event with significant confidence in low-latency (IFAR higher than 1 year), recovered it in their offline analyses. These searches differ from PyCBC (Davies:2020tsx; Chandra:2021wbw; Davis:2022cmw; Kumar:2024bfe) in their implementation and use of signal–noise discriminators. GstLAL’s enhanced significance (the higher IFAR in Table. 1) is primarily driven by a higher mass extension of the search with specific settings to compute the background for such higher mass mergers accurately. More details of the settings are provided in (GWTC:Methods; Joshi:2025nty). These changes improved the signal and noise models in this part of the parameter space in general, leading to a better recovery of this high mass signal. Additionally, cWB-GMM, an entirely offline model-independent search, uses Gaussian Mixture Models (Gayathri:2020bly; Lopez:2021ikt; Smith:2024bsn) to rerank the triggers identified by cWB-2G.
@@
The differences in the IFARs reported by the offline search pipelines—despite broadly consistent signal-to-noise ratios—primarily reflect differences in their ability to separate GW231123-like signals from background noise in a comparable parameter range. Similar discrepancies have been observed previously, particularly between matched-filter and minimally-modelled searches, when searching for non-eccentric intermediate-mass black hole (IMBH) binaries (CalderonBustillo:2017skv; Chandra:2020ccy; LIGOScientific:2020iuh; Chandra:2021xvs; Szczepanczyk:2020osv). These differences arise not only from how effectively each search separates signals from glitches, but also from the differing approaches used to estimate the noise background.
To assess whether the observed variation in statistical significance across pipelines is consistent with expectations, we conducted a dedicated injection campaign. Using the NRSur7dq4 (NRSur) waveform model (Varma:2019csw), we simulated 8000 non-eccentric BBH signals with intrinsic parameters consistent with those inferred for GW231123 (Section 4). We sampled the sky positions and binary orientations isotropically and drew redshifts uniformly in comoving volume up to , assuming a flat CDM cosmology (Planck:2015fie). We added these simulated signals uniformly over several days around the event and re-ran our offline search pipelines using the same configuration as applied to the real data.
@@
Given that all pipelines detected GW231123 with an IFAR above the typical threshold of 1 year used for population analyses, and a detailed background study for one pipeline (cWB-BBH) identified GW231123 with an IFAR of 9700 years, we consider GW231123 to be a confident detection.
@@ 3 Data quality
The event GW231123 was detected during the first part of the fourth observing run (O4a), a time when the LIGO Hanford and LIGO Livingston detectors were observing with a typical binary neutron star inspiral range of 152 Mpc and 160 Mpc (Capote:2024rmo). The detectors’ data were calibrated in near real-time to produce the online dataset used for low-latency searches (abbott2020guide; klimenko2016method; Tsukada:2023edh; Ewing:2023qqe; DalCanton:2020vpm; Chu:2020pjv; Aubin:2020goo) and parameter estimation (Singer:2015ema; bilby_paper; Pankow:2015cra). The calibration process subtracts linear spectral features from known instrumental sources, identified through auxiliary witness sensors, and intentionally injected calibration lines used to measure the instruments’ response at various frequencies (viets2018reconstructing; Sun:2020wke; Sun:2021qcg).
@@
Following the data-quality procedures established for O4a (LIGO:2024kkz), including broadband noise subtraction (Vajente:2019ycy) and data-quality report analysis routines (LIGO:2021ppb), detector data surrounding the event were evaluated for signs of non-Gaussian excess power (glitches) within the target time-frequency analysis window using a spectrogram-based glitch-identification tool (Vazsonyi:2022jul). It was determined that glitches were present in each detector around, but not coincident with the event.
@@
From spectrograms, we determined that a glitch was present in the LIGO Hanford data 1.7–1.1 s before the event, in a frequency range between 15–30 Hz. The glitch is possibly related to the LIGO Hanford differential arm control loop (LIGOScientific:2014pky). This control loop leads to nonstationary noise from the high root-mean-square drive applied to the electrostatic drive actuator. This issue has been fixed in the second part of the fourth observing run (darm_improvement). This glitch was close to the event and within the time–frequency window used to infer the source properties, so BayesWave (Cornish:2014kda; Cornish:2020dwh; Chatziioannou:2021ezd) was used to model simultaneously the compact binary signal and the glitch (LIGO:2024kkz). We removed this non-Gaussianity from the data by subtracting a phenomenological, wavelet-based model of the excess power noise (Hourihane:2022doe; Ghonge:2023ksb). The glitch-subtracted data successfully passed the validation process, which compares the residual noise to Gaussian noise (LIGO:2024kkz; Vazsonyi:2022jul). Additional broadband non-stationary noise was present in the Hanford detector in the hours of data surrounding GW231123, but we found no evidence that this impacted the analysis of GW231123.
@@
In LIGO Livingston data, a glitch was identified 3.0–2.0 s before the event, in a frequency range between 10–20 Hz. Given that LIGO Livingston had recurring low-frequency scattered light glitches (LIGO:2024kkz), this glitch was likely caused by scattered light. We determined the time–frequency profile of the glitch to have no measurable effect on the GW231123 analysis, so the analyses from here on use the LIGO Livingston original data and the LIGO Hanford glitch-subtracted data.
@@ 4.1 Methods
We report the properties of GW231123 using signal models for non-eccentric BBH mergers in a coherent Bayesian analysis (LIGOScientific:2016vlm) of the LIGO Hanford and LIGO Livingston data around the time of GW231123. (We discuss potential eccentricity further in Section 6.) We calculate the likelihood using 8 s of data (6 s before and 2 s after the reported merger time of GW231123), and consider frequencies within the range 20–448 Hz. This range was chosen to contain the signal based on preliminary analyses at the time of the event, and to avoid loss of power at high frequencies due to low-pass filtering of the data (LIGOScientific:2025yae). All analyses employ standard priors used in previous analyses (LIGOScientific:2020ibl; LIGOScientific:2021usb; LIGOScientific:2021vkt), and we use a Planck 2015 cosmology (Planck:2015fie). The NRSur analysis employs a reduced prior mass range due to model constraints, mass ratios below 6:1. The other models employ a wider mass prior, mass ratios below 10:1, and no posterior support is found beyond the NRSur analysis. We characterise the detector noise via the median estimate of different power spectral density (PSD) realizations calculated with BayesWave (Cornish:2014kda; Littenberg:2014oda). As done previously (Chatziioannou:2019zvs; LIGOScientific:2021vkt), we calculate the median PSD for data containing the trigger. To sample the posterior distribution, we interface with the dynesty nested sampling package (Speagle:2019ivv) via the bilby library (bilby_paper; Romero-Shaw:2020owr). We present results with 1000 live points, and verify that the results remain consistent when the number of live points is increased to 3000, as well as when we lower the frequency range to include data between 16–20 Hz.
@@ 4.2 Waveform Systematics
The source properties of GW231123 lie in a challenging region of parameter space for current waveform models, to such an extent that measurements using different models show significant disagreement, with multiple parameters failing to agree within 90% credible intervals. (See Appendix A for examples.) For typical signals, our models are well within our observations’ accuracy requirements, and the level of model disagreement for GW231123 has not been seen in any previous LVK GW observation with moderate SNRs (12). All models show strong support for spins 0.8, and since no theoretical signal model is calibrated to numerical-relativity (NR) waveforms from precessing binaries with spins above 0.8, waveform uncertainties are one possible cause of the measurement differences. Hence, before presenting the source properties, we describe how we quantify waveform-model uncertainties. We do not study in detail the impact of Gaussian noise fluctuations or low-SNR glitches that are difficult to identify and mitigate using the methods presented in Section 3.
@@
We consider five state-of-the-art inspiral–merger–ringdown (IMR) signal models, NRSur7dq4 (NRSur; Varma:2019csw), SEOBNRv5PHM (v5PHM; Ramos-Buades:2023ehm), IMRPhenomTPHM (TPHM; Estelles:2021gvs), IMRPhenomXPHM (XPHM; Colleoni:2024knd) and IMRPhenomXO4a (XO4a; Thompson:2023ase). The first three model the signal in the time domain while the latter two natively employ the frequency domain. All models use information from numerical relativity to inform the merger-ringdown in the aligned-spin sector. However only NRSur; fully interpolates two-spin precessing systems in the precessing sector, while XO4a is calibrated to single-spin precessing systems; all other models employ results from post-Newtonian and perturbation theory through merger and ringdown. (More details are given in Appendix A.) The model papers referenced here include studies to assess the accuracy of these models across the BBH parameter space, but here we focus on the likely region of parameter space for this observation; high total mass, , and moderate to high spins.
@@
We quantify the models’ accuracy against NR results, including a set of simulations that extend up to spins of 0.95 (Boyle:2019kee; Hamilton:2023qkv; Scheel:2025jct). A standard waveform accuracy measure is the mismatch between two waveforms (Cutler:1994ys), where waveform uncertainties will not bias a parameter measurement if the model’s mismatch uncertainty is less than (McWilliams:2010eq; Baird:2012cu), where is the SNR and is the chi-square value for degrees of freedom at probability . For single-parameter measurements provides a lower bound (Thompson:2025ab), so the mismatch criterion for the 90% credible interval at is . Figure 2 reports the distribution of mismatches of each model against 1123 NR waveforms with , all scaled to the redshifted (detector-frame) total mass , at six equally spaced inclinations in from to inclusive. The mismatches are calculated for precessing systems (Schmidt:2014iyl; Harry:2016ijz) following the procedure described in Hamilton:2021pkf, maximising over time shifts, a global phase and template polarisation and optimising over in-plane spin rotations. A subset of the simulations come from the third release of the SXS catalog (Scheel:2025jct) and contain GW memory, which introduces a constant late-time offset that we handle for the mismatch calculations with a highpass filtering technique (Xu:2024ybt; Valencia:2024zhi; Chen:2024ieh) to mitigate possible artefacts in the Fourier domain. We employ the same PSD for the LIGO Livingston detector as utilised in the coherent Bayesian analysis. NRSur performs better than the other models (by roughly an order of magnitude for low-spin cases), and all other models have comparable accuracy. However, NRSur does not meet the conservative accuracy criterion for all cases, and for spins greater than 0.8 the mismatches are higher in 10% of 98 cases. Even if we apply a less conservative mismatch criterion from the literature (e.g., with for non-eccentric binaries and (Chatziioannou:2017tdw; Scheel:2025jct) we have 0.0072) there are configurations where NRSur exceeds the criterion (2%). The relative accuracy of models is also not uniform across all cases, e.g., we find cases in which other models show comparable or improved performance relative to NRSur.
@@
To test whether these waveform uncertainties will result in biases, we performed our standard Bayesian parameter estimation analysis on a series of NR injections, as detailed in Appendix A. We observe that, while the five models considered here perform well for most signals, there are configurations where all models may incur biases for massive high-spin signals. We also find that the relative performance of each model can change in the presence of Gaussian noise, although this requires more detailed study in future work. To properly correct for this, we would ideally marginalise over waveform uncertainties or incorporate model accuracy into Bayesian analyses (Read:2023hkv; Khan:2024whs; Hoy:2024vpc; Pompili:2024yec; Kumar:2025nwb; Mezzasoma:2025moh). Without access to a model of the waveform-model uncertainties, we follow what has been done previously (LIGOScientific:2016vlm) and combine the results from multiple models to marginalise over the model uncertainties. In choosing models in addition to NRSur, we note that all other models exhibit a comparable range of mismatches, and no model is clearly preferred in our injection studies, and so we include all five state-of-the-art models. We combine posterior results inferred from NRSur, v5PHM, TPHM, XPHM, XO4a with equal weight and report the combined samples throughout this paper. To illustrate the variation between the combined results and single models, in some figures we also show the NRSur results. In some analyses we expect the choice of model to have little impact, e.g., the detection significance study in Section 2, and in these cases, we use only the NRSur samples.